# Introspective Growth: Automatically Advancing LLM Expertise in Technology Judgment

### v2

@@ Appendix B Related Literature
Self-questioning and self-growth of LLMs: LLMs have recently been prompted to engage in self-questioning and introspection to improve their reasoning and conceptual understanding [25, 46, 51, 61, 60]. Instead of only generating direct answers, many contemporary models are encouraged to ask themselves clarifying questions and work through sub-problems, mimicking a Socratic or pedagogical approach. Early prompting techniques like chain-of-thought (CoT) [62] demonstrated that guiding LLMs to reason step-by-step markedly boosts performance on complex tasks. Building on this, researchers introduced methods for the model to generate its own intermediate questions. For example, self-ask prompting [43] explicitly guides the model to pose follow-up questions to itself and then answer them in sequence, effectively decomposing complex queries into tractable pieces. Beyond decomposition, recursive self-reflection techniques have emerged to let LLMs iteratively improve their answers. Madaan et al. [34] introduce Self-Refine, a framework where an LLM generates an initial answer and then critiques and revises its own output in multiple rounds. Shinn et al. [48] propose Reflexion, in which an agent-model “verbally reflects” on feedback from previous trials and stores these self-generated insights in memory to inform subsequent attempts. Huang et al. [25] showed that even without external labels, an LLM can “self-train” by generating reasoning traces for unsolved problems and fine-tuning on them, yielding notable gains in mathematical and commonsense tasks. Miao et al. present SelfCheck [37], a zero-shot verification schema enabling LLMs to assess their own step-by-step reasoning without external resources, thereby enhancing accuracy in complex tasks through internal consistency checks. Similarly, techniques like least-to-most prompting [65] encourage the model to teach itself gradually, which has improved performance on multi-step reasoning challenges. Other work has explored letting models self-evaluate or re-rank their answers (e.g., self-consistency voting [59]) and prompts them to explicitly verify or justify their solutions, analogous to a student checking their work. These introspective or pedagogically-inspired strategies are motivated by human learning: students often learn better by asking themselves questions and reflecting on mistakes. Empirically, such approaches can bolster an LLM’s conceptual grasp and reduce reasoning errors. Nevertheless, there are also limitations and mixed findings. It is hard to evaluate the quality of self-generated questions because LLMs may not be able to answer their self-generated questions [5]. As we demonstrate, large models are much more likely to productively answer small model questions than the converse. Not all tasks benefit uniformly from self-reflection. Li et al. [33] find that prompting GPT-4 to reflect on its answers helps on a truth-focused QA task but hinders performance on a multi-hop knowledge task. Thus, while self-questioning and introspective prompts represent a powerful paradigm for enhancing LLM reasoning, careful tuning is needed to ensure the model’s “inner dialogue” remains productive and does not overthink or hallucinate.
@@ Appendix E Prompts
Few-shot CoT I give you two examples: Example 1 Abstract 1: A method is provided for transmitting control information between a line-switching and a packet-switching communications network. A signaling connection is set up in a packet-switching network in order to transmit signaling packets which form connection-independent control information which relates to at least one service feature in a line-switching communications network, so that at least one service feature of the line-switching communications network may be used in the packet-switching communications network by means of the control information independently of the connection. Abstract 2: A method and apparatus for connecting networks of different types of transmission wherein, in a first network, the data are transmitted in a circuit-switched manner, and in a second network, the data are packet-switched and transmitted connectionlessly on the lower three protocol levels. To be able to perform the signaling in a simple manner, signaling units are used which process signaling messages of the circuit-switched network. Label: The two abstracts are describing different patents according to human expert judgment. Example 2 Abstract 1: A method is provided for transmitting control information between a line-switching and a packet-switching communications network. A signaling connection is set up in a packet-switching network in order to transmit signaling packets which form connection-independent control information which relates to at least one service feature in a line-switching communications network, so that at least one service feature of the line-switching communications network may be used in the packet-switching communications network by means of the control information independently of the connection. Abstract 2: A method is disclosed for transmitting control information between a line-switching and a packet-switching communications network. A signaling connection is established within the packet-switching network to transmit signaling packets that carry connection-independent control information. This control information pertains to at least one service feature of the line-switching communications network, enabling the use of that service feature within the packet-switching communications network independently of any specific connection. Label: The two abstracts describe the same patent, as the second is a rephrasing of the first. // Then attach the Prompt for the Splitting Task Judgment
Appendix F External Knowledge and Compressed Knowledge
In the original paper, we identify a consistent gap between the internal knowledge of LLMs and external scientific knowledge (see Figure 5 in the original paper). We propose two possible explanations for this gap: (1) LLMs function as compressed representations of knowledge. Scientific and technological information, in particular, tends to be sparsely and unevenly distributed in the training corpus, and the compression process significantly degrades the fidelity of such underrepresented knowledge. (2) LLMs acquire new knowledge at inference time — that is, they attempt to reason about or generate responses related to concepts they have never encountered during training but acquired in scientific retrieval.
Both mechanisms may contribute to the observed gap, we argue that the first is the primary cause, for the following reasons:
(a) Panel a
(b) Panel b
Figure 9: Panel a: sample questions. Panel b: scientific retrieval does not introduce more "content-level" new information for LLM to answer these questions.
  • Models tested in the paper — such as LLaMA 3.1, released in July 2024777https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, when approximately 81% of the papers that produced scientific answers had already been posted, and LLaMA 3.2, released around October 2024888https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct, when approximately 86% of such papers were available — should have had access to most relevant information.
  • Qualitative inspection of the generated questions suggests that answering them does not require substantial novel information beyond the models’ training. Many questions are straightforward and focus on established concepts or systems. For example, questions involving the TCP Port Service Multiplexer (TCPMUX) relate to long-standing infrastructure, with patents contributing only incremental advancements. As shown in our original paper, 68% of the Llama 1B model’s questions and 44% of the Llama 405B model’s questions include phrases like “what is” or “what are,” indicating the dominance of factual, definitional queries. A sample of questions can be seen in Figure 9 panel a.
  • We further analyzed this by embedding all answers from the scientific sources and the LLM’s (LLaMA 8b) self-talk using the model from [15], specifically pretrained via contrastive learning on patent citation relationships. To visualize and compare the distribution of embedding spaces, we reduce the dimension to 2, estimate the boundary of each region using 5% Gaussian kernel density estimation — that is, we draw contours enclosing the densest 95% of the data points in each embedding set. If LLMs substantially incorporate new scientific content, we would expect their embedding regions to diverge meaningfully from those derived directly from their self-talk. However, as shown in Figure 9 panel b, the embedding regions largely overlap. Interestingly, self-talks display a distinct outlier region, which we suspect may reflect hallucinated content introduced by the LLM.
Despite the similar content - the shared 2D embedding space that represents the greatest variance in high (768)-dimension space, these two types of answers still differ in several meaningful ways. Scientific answers, averaging 62.1$\pm$1.8 words, are possibly richer in detail, drawing more heavily from scientific evidence. In contrast, self-talk answers are notably shorter, averaging 38.2$\pm$0.6 words — only half the length of scientific responses.
Figure 10: The permutation test proves that two types of answers differ significantly in details.
We conducted a permutation test on the centroid distance between the two groups to assess whether self-talk and scientific answers originate from different distributions in embedding space (Figure 10). The test evaluates the null hypothesis that the two groups are exchangeable—that is, their embeddings are drawn from the same underlying distribution. We first computed the Euclidean distance between the centroids of the two groups, which was 1.16 (red line). Then, we randomly permuted the group labels 1,000 times and recalculated the centroid distance for each permutation to generate a null distribution. We find that the observed separation is unlikely under the null hypothesis (reject null, p=0.000). This provides strong statistical evidence that self-talk and scientific answers, although they overlap in coarse-grained content, differ significantly in detail.
Appendix G Results for Majority Voting Schema
@@
G.1 Performances of other LLMs with varying sizes on splitting patents
@@
G.2 Scientific QAs benefit LLMs’ understanding
(a) QA from science overperforms same questions + placebo answers.
(b) Scientific QA reduces confidence in wrong answers.
(c) Scientific QA reduces wrong answers’ consistency, raises correct answers’ consistency.
Figure 14: Ask Llama 3.1 8B instruct "Do you think they are the same patent?" (splitting task). 1 = yes is the wrong answer.
G.3 LLMs do not know what they have known
@@
G.4 Self-talk QAs VS. Chain-of-Thought
Figure 16: Self-talk QAs and CoT prompting manifest different performances.
Figure 17: Self-talk QAs and Few-shot CoT prompting converge.
G.5 Smaller models’ questions help mid-size models more than those from larger models
@@
Appendix H Discussion of Limitations and Future Work
Several limitations and directions for future work remain. First, we note that certain patent pairs could be effectively indistinguishable based on abstract-level information alone. In such cases, even exhaustive self-questioning by big models fails to resolve ambiguity, suggesting an upper bound on performance determined by input granularity. This highlights the need for richer representations (e.g., claims, diagrams, or specification sections) in future benchmarks. Another, even more intriguing, possibility is that LLMs give the “same” label because they sometimes perceive the underlying similarities between genuinely similar, and potentially follow-on or copying patents. Meanwhile, human judges can also be misled by the storytelling of patent applicants emphasizing nonexistent novelty. In other words, recognizing similarities rather than differences may reflect a more advanced capability for LLMs — a possibility that warrants further investigation.
Second, although our current setup uses scientific retrieval from arXiv papers, a more promising direction is to evolve the retrieval process itself. Rather than relying on fixed sources or searching questions from scientific databases, LLMs could learn to refine or even generate better retrieval queries via reinforcement learning, active learning, or model-in-the-loop optimization based on science or what they have learned from prior reasoning steps. An agentic framing - where models, like a baby sensing the world, autonomously refine their questions and knowledge, decide whether to retrieve information, evaluate whether to accept or reject what is retrieved, and determine how much knowledge to seek for one question/how many questions to ask - could enable deeper forms of exploration beyond single-pass self-questioning.
Third, the exact nature of “resolution” in distinguishing patents remains underdefined. That is, even when models succeed at effectively splitting, it is not always clear what semantic threshold they are using. Are they resolving based on functionality, application, architecture, or terminology? To what extent are they leveraging scientific knowledge—which aspects, and how relevant is that knowledge to the task? Future work could explicitly model and evaluate these questions to understand what kinds of conceptual distinctions LLMs are truly capable of making by probing their internal, layer-wise conceptual representations and dynamics. It will also be interesting to look into differences in judgment performance across patent classes and categories, as LLM internal knowledge likely varies across fields.
Appendix I Ethics Statement
This research adheres fully to the Code of Ethics. The work does not involve human subjects, crowdsourcing, or sensitive or private data. It does not pose foreseeable risks of misuse, and therefore, no special safeguards are necessary. All external assets, including datasets and pretrained models, are properly credited and used in accordance with their respective licenses. Any new datasets introduced are derived from publicly available sources, anonymized as appropriate, and documented to ensure transparency and reproducibility. Societal impact has been briefly discussed, and the study does not require specific mitigation strategies.

### v3

@@ Appendix B Related Literature
Self-questioning and self-growth of LLMs: LLMs have recently been prompted to engage in self-questioning and introspection to improve their reasoning and conceptual understanding [Huang et al., 2023, Qu et al., 2024, Song et al., 2024, Wang et al., 2023c, b]. Instead of only generating direct answers, many contemporary models are encouraged to ask themselves clarifying questions and work through sub-problems, mimicking a Socratic or pedagogical approach. Early prompting techniques like chain-of-thought (CoT) [Wei et al., 2022] demonstrated that guiding LLMs to reason step-by-step markedly boosts performance on complex tasks. Building on this, researchers introduced methods for the model to generate its own intermediate questions. For example, self-ask prompting [Press et al., 2022] explicitly guides the model to pose follow-up questions to itself and then answer them in sequence, effectively decomposing complex queries into tractable pieces. Beyond decomposition, recursive self-reflection techniques have emerged to let LLMs iteratively improve their answers. Madaan et al. [Madaan et al., 2023] introduce Self-Refine, a framework where an LLM generates an initial answer and then critiques and revises its own output in multiple rounds. Shinn et al. [Shinn et al., 2023] propose Reflexion, in which an agent-model “verbally reflects” on feedback from previous trials and stores these self-generated insights in memory to inform subsequent attempts. Huang et al. [Huang et al., 2023] showed that even without external labels, an LLM can “self-train” by generating reasoning traces for unsolved problems and fine-tuning on them, yielding notable gains in mathematical and commonsense tasks. Miao et al. present SelfCheck [Miao et al., 2024], a zero-shot verification schema enabling LLMs to assess their own step-by-step reasoning without external resources, thereby enhancing accuracy in complex tasks through internal consistency checks. Similarly, techniques like least-to-most prompting [Zhou et al., 2022] encourage the model to teach itself gradually, which has improved performance on multi-step reasoning challenges. Other work has explored letting models self-evaluate or re-rank their answers (e.g., self-consistency voting [Wang et al., 2023a]) and prompts them to explicitly verify or justify their solutions, analogous to a student checking their work. These introspective or pedagogically-inspired strategies are motivated by human learning: students often learn better by asking themselves questions and reflecting on mistakes. Empirically, such approaches can bolster an LLM’s conceptual grasp and reduce reasoning errors. Nevertheless, there are also limitations and mixed findings. It is hard to evaluate the quality of self-generated questions because LLMs may not be able to answer their self-generated questions [Balepur et al., 2025]. As we demonstrate, large models are much more likely to productively answer small model questions than the converse. Not all tasks benefit uniformly from self-reflection. Li et al. [Li et al., 2024] find that prompting GPT-4 to reflect on its answers helps on a truth-focused QA task but hinders performance on a multi-hop knowledge task. Thus, while self-questioning and introspective prompts represent a powerful paradigm for enhancing LLM reasoning, careful tuning is needed to ensure the model’s “inner dialogue” remains productive and does not overthink or hallucinate.
@@ Appendix E Prompts
Appendix F Results for Majority Voting Schema
@@
F.1 Performances of other LLMs with varying sizes on splitting patents
@@
F.2 LLMs do not know what they have known
@@
F.3 Smaller models’ questions help mid-size models more than those from larger models
@@
Appendix G Discussion of Limitations and Future Work
Several limitations and directions for future work remain.
Although our current setup uses scientific retrieval from arXiv papers, a more promising direction is to evolve the retrieval process itself. Rather than relying on fixed sources or searching questions from scientific databases, LLMs could learn to refine or even generate better retrieval queries via reinforcement learning, active learning, or model-in-the-loop optimization based on science or what they have learned from prior reasoning steps. An agentic framing - where models, like a baby sensing the world, autonomously refine their questions and knowledge, decide whether to retrieve information, evaluate whether to accept or reject what is retrieved, and determine how much knowledge to seek for one question/how many questions to ask - could enable deeper forms of exploration beyond single-pass self-questioning.
The exact nature of “resolution” in distinguishing patents remains underdefined. That is, even when models succeed at effectively splitting, it is not always clear what semantic threshold they are using. Are they resolving based on functionality, application, architecture, or terminology? Future work could explicitly model and evaluate these questions to understand what kinds of conceptual distinctions LLMs are truly capable of making by probing their internal, layer-wise conceptual representations and dynamics. It will also be interesting to look into differences in judgment performance across patent classes and categories, as LLM internal knowledge likely varies across fields.